Can you help me with these i want to make sure i did it all right
(x-5)^2+2(x-5)-35=0
(x-2)^2-3(x-2)+2=0
2(x^2-5)^2-13(x^2-5)+20=0
2x(2/3)+5(1/3)=12
6X(2/3)-5X(1/3)-6=0
X^-2+4X^-1=12
X^4-8X^2+7=0
X^4-2X^2-35=0

Answer by Alan3354(69443)

That's too many problems.
Repost with your answers, and I'll check them for you.
---------
I'll do one.
(x-5)^2+2(x-5)-35=0
Sub y for x-5
y^2 + 2y - 35 = 0
(y+7)*(y-5) = 0
y = -7, 5
-----------
x-5 = -7 --> x = -2
x-5 = 5 --> x = 10